# Adaptive Nonlinear Model Predictive Control for a Real-World Labyrinth Game

### v1

@@ III-A State-Space Model
Considering only the ball-plate system without obstacles or friction, the dynamics are given by [10] as
@@ III-C Model Predictive Control
A typical situation from the labyrinth is shown in Fig. 5. The task of our high-level path-planning MPC (HL-solver) is to find or approximate the optimal path around the obstacles (purple). Based on the current ball position, we choose the third next corner on the underlying perpendicular path (grey) as the current goal (green star) for the HL-solver and update it in a receding horizon scheme. A more distant goal could result in more optimal paths but makes the problem harder to solve. We incorporate the ball-plate system dynamics and set the obstacles as constraints. The solution that the HL-solver finds is a locally optimal path (red), which approximates the globally optimal path. To track the path, we propose a low-level MPC (LL-solver) that computes a control input for each new state estimate at 55Hz. It incorporates the ball-plate-system dynamics but is not constrained by the obstacles, which reduces the solving time drastically. Splitting the system in a slow HL-solver and a fast LL-solver allows us to apply the optimal solutions of computationally heavy MPC in a high-frequency real-time environment. An overview can be seen in Table I.
@@ III-C1 High-Level Solver
The objective function’s focus lies on reaching the goal while ensuring stable behavior through small velocity and input weights
@@
We constrain state and input by the labyrinth frame () and reduced values of the maximal possible angles () and servo velocity . The initial and final constraints are set to the current state estimate and the current goal () (3rd next corner). The inequality constraints are formulated as a stack of nonlinear functions with fixed upper and lower bounds. They incorporate the obstacles, i.e., the walls and holes. The distance of the ball center to the hole center (see Fig. 6) is
In order that the ball does not fall into the hole we constrain . We reformulate to the form of equation 15
@@
In every controller iteration, a look-up table is used to only choose the closest 5 holes and 10 walls to incorporate in the optimization problem. Obstacle positions and size are given to as run-time parameters . Especially for non-convex problems, the initial guess for the solver is crucial for convergence to the right solution. According to choosing the goal, we use the second next corner of the labyrinth () as the initial guess
The HL-solver yields a feasible 100-step path including ball position, velocity, angles, and system inputs.
@@ IV Results
For comparison, we introduce a cascaded PID controller and a linear MPC contoller. Both controllers follow a trajectory of given way-points. Once the ball comes close to a way-point (7mm), the next way-point is chosen. The PID’s cascaded structure is shown in Fig. 10. The linear MPC uses the same state-space model from Equation 8 as its interstage-equality constraint. Its objective function minimizes the distance to the next way-point and penalizes ball velocity and input with small weights. Information about the obstacles is only incorporated in a way that one way point is always set so that the next one is in line of sight. Furthermore, the disturbance compensator (Section III-C3) is applied as well.
We perform 25 runs for each controller and measure the achieved distance by means of noting the hole number where the ball fell in, and record the completion time for successful runs. These results are summarized in Table II. As can be seen, both MPC approaches are able to do full runs on the labyrinth. Nevertheless, the nonlinear, non-convex approach incorporating the obstacles in its planning achieves a higher success rate with lower completion times. The PID shows the correct behavior but is not able to complete the full labyrinth.
@@ V Conclusion
We presented an approach of splitting a computation-heavy MPC problem for a real-world labyrinth game into two sub-problems, to leverage the benefits of formulating a nonlinear non-convex optimization problem and still applying it to a high-frequency real-time system. We introduced a method to incorporate the differently shaped obstacles in the optimization problem and obtain a competitive controller that can deal with a variety of disturbances. The resulting controller is able to solve the labyrinth without any mechanical modifications.
In future work, we envision increasing the overall robustness of the controller in two ways. First, a more accurate model can be obtained by modeling the dynamics of the plate tilting mechanism. Second, modeling the collision of the ball with the walls could lead to more robust strategies where the collision is exploited to redirect the ball, instead of attempting to avoid the walls.

### v2

@@ III-A State-Space Model
Considering only the ball-plate system without obstacles or friction, the dynamics for a solid ball are given by [10] as
@@ III-C Model Predictive Control
A typical situation from the labyrinth is shown in Fig. 5. The task of our high-level path-planning MPC (HL-solver) is to find or approximate the optimal path around the obstacles (purple). Based on the current ball position, we choose the third next corner on the underlying perpendicular path (grey) as the current goal (green star) for the HL-solver and update it in a receding horizon scheme. A more distant goal could result in more optimal paths but makes the problem harder to solve. The corner locations of the underlying perpendicular path are given to the system. We incorporate the ball-plate system dynamics and set the obstacles as constraints. The solution that the HL-solver finds is a locally optimal path (red), which approximates the globally optimal path. To track the path, we propose a low-level MPC (LL-solver) that computes a control input for each new state estimate at 55Hz. It incorporates the ball-plate-system dynamics but is not constrained by the obstacles, which reduces the solving time drastically. Splitting the system in a slow HL-solver and a fast LL-solver allows us to apply the optimal solutions of computationally heavy MPC in a high-frequency real-time environment. An overview can be seen in Table I. The unit of length in the problem formulations is m, angles in rad, time in s and the weights are unitless.
@@ III-C1 High-Level Solver
The objective function’s focus lies on reaching the goal while ensuring smooth behavior through small velocity and input weights
@@
We constrain the state and input using the labyrinth frame () and reduced values of the maximum possible angles () and servo velocity . The initial and final constraints are set to the current state estimate and the current goal () (3rd next corner). The inequality constraints are formulated as a stack of nonlinear functions with fixed upper and lower bounds. They incorporate the obstacles, i.e., the walls and holes. The distance of the ball center to the hole center (see Fig. 6) is
In order that the ball does not fall into the hole we constrain . We reformulate to the form of Equation (15)
@@
To reduce computation time, a look-up table is used to only choose the closest 5 holes and 10 walls to the ball to incorporate in the optimization problem. Obstacle positions and size are given to as run-time parameters . Especially for non-convex problems, the initial guess for the solver is crucial for convergence to the right solution. According to choosing the goal, we use the second next corner of the labyrinth () as the initial guess
The HL-solver yields a feasible 100-step path including ball position, velocity, angles, and system inputs. In case the solver doesn’t find a solution for 200 ms, we terminate and restart the solver from the new state. Until the next solution is found, we use the last feasible HL-solution.
@@ IV Results
For comparison, we introduce a cascaded PID controller and a linear MPC contoller. Both controllers follow a trajectory of given way-points. Once the ball comes close to a way-point (7mm), the next way-point is chosen. The PID’s cascaded structure is shown in Fig. 10. The linear MPC uses the same state-space model from Equation 8 as its interstage-equality constraint. Its objective function minimizes the distance to the next way-point and penalizes ball velocity and input with small weights. Information about the obstacles is only incorporated in a way that one way point is always set so that the next one is in line of sight. The disturbance compensator (Section III-C3) and the feed-forward angle map (Section III-C4) are applied as well.
We perform 25 runs for each controller and measure the achieved distance by means of noting the hole number where the ball fell in, and record the completion time for successful runs. These results are summarized in Table II and depicted in Fig. 11. As can be seen, both MPC approaches are able to do full runs on the labyrinth. Nevertheless, the nonlinear, non-convex approach incorporating the obstacles in its planning achieves a higher success rate with lower completion times. The PID shows the correct behavior but is not able to complete the full labyrinth.
@@ V Conclusion
We presented an approach of splitting a computation-heavy MPC problem for a real-world labyrinth game into two sub-problems, to leverage the benefits of formulating a nonlinear non-convex optimization problem and still applying it to a high-frequency real-time system. We introduced a method to incorporate the differently shaped obstacles in the optimization problem and obtain a competitive controller that can deal with a variety of disturbances. The resulting controller is able to solve the labyrinth without any mechanical modifications. We envision that our method can be applied to similar control problems e.g. autonomous race cars on an obstacle course [14]. In future work, we envision increasing the overall robustness of the controller in two ways. First, a more accurate model can be obtained by modeling the dynamics of the plate tilting mechanism including the spring. Second, modeling the collision of the ball with the walls could lead to more robust strategies where the collision is exploited to redirect the ball, instead of attempting to avoid the walls.